# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v1

@@ IV-B4 Automatic Program Repair
Buggy code can take a lot of effort to fix. Automatic program repair leverages generative models to output the correct version. RAG techinque is widely used for few-shot learning in automatic program repair [190, 130, 131, 174, 189]. Among them, RING [190] retrieves similar error messages based on both sparse and dense vectors, then builds prompt for the generator Codex [2]. CEDAR [130] applies for both assertion generation and program repairs tasks; it uses sparse and dense retrieval to search for similar codes, then forms prompt for Codex to generate results. InferFix [131] crafts a prompt carrying the bug type, location, relevant syntax hierarchies, and similar fixes through dense retrieval. Then it also uses Codex for generation. RAP-Gen [174] also retrieves similar buggy codes and corresponding fixes through hybrid retriever (including both sparse and dense retriever). It finetunes CodeT5 [105] with this RAG paradigm. SARGAM [189] searches similar buggy code using dense retrieval, generates patches using augmented input, then applies another model to modify the final result. These research works also involve error localization, which is not our focus.
@@ IV-B5 Text-to-SQL and Code-based Semantic Parsing
Semantic parsing is the task of translating natural language utterances to unambiguous structured meaning representations, where code language is often leveraged to augment this process. SQL is not only a programming language but can also be considered as a structured representation, so we place text-to-SQL (a special case of code generation) in this subsection. Related research works all apply retrieval-augmented few-shot learning. XRICL [178] focuses on the problem of translating non-English utterances into SQL queries. It searches and reranks English utterance using non-English ones by dense retrieval, then builds prompt for Codex to generate SQL queries. SYNCHROMESH [120] proposes constrained semantic decoding to enforce rich syntactic and semantic constraints during generation of SQL or other languages. It uses the similarity between abstract syntax trees (AST) to finetune the dense retriever S-Bert. During inference, similar NL and SQL are retrived to form the prompt of GPT-3. CodeICL [225] uses Python for semantic parsing, and augments prompts with a structured domain description for GPT generation. In few-shot learning setting, it leverages BM25 sparse retrieval to find similar training examples.
@@ IV-B6 Others
There are several other code-related tasks that adopt RAG paradigm. In  [226] for numerical reasoning task, the Chain-of-Thought is replaced with the programs as the intermediate reasoning step, and dense retrieval-based similar examples are augmented in prompt for downstream LLMs. E&V [227] leverages an LLM-based agent for program static analysis. The agent uses source code retrieval, pseudo-code execution, execution specifications verification, and other tools to form intermediate results. The retrieval is implemented by AST pattern matching. StackSpotAI [228] builds an AI coding assistant, which incorporates many code-related tasks. It implements an RAG component, identifying the most relevant pieces of information which serve as the context for GPT generation.
@@ IV-C1 Audio Generation
The goal of audio generation is to generate audio with natural language input. Re-AudioLDM [157] adopts dense retriever CLAP [25] to retrieve similar caption-audio pairs given input prompt. The generator, latent diffusion model and VAE-based decoder, take the representations of input text and retrieved pairs as input and generate output audio. Make-An-Audio [42] uses dense retriever CLAP [25] to augment data, retrieving related audio given natural language text. It then constructs pseudo prompts for diffusion-based text-to-audio model training.
@@ IV-D1 Image Generation
RAG enhances generative models by incorporating an information retrieval system. For image generation [43, 137, 149, 150, 150, 148, 233, 151, 152], this means leveraging retrieved data to produce high-fidelity and faithful images, even for rare or unseen entities. Meanwhile, this approach can also reduce the parameter count and computational cost of the generative model.
RetrieveGAN [43] uses a differentiable retrieval process to select compatible patches from other images as reference for the generation. It adopts the Gumbel-softmax trick to make the retrieval process differentiable, enabling the end-to-end training and the optimization of the embedding function. It also encourages the selection of mutually compatible patches with additional objective functions. RetrieveGAN can generate realistic and diverse images from scene descriptions, where the retrieved patches are reasonable and coherent. IC-GAN [137] models the data distribution as a mixture of conditional distributions around each training instance. It conditions both the generator and the discriminator on instance features, which are obtained from a pre-trained feature extractor. It also uses the nearest neighbors of the conditioning instance as real samples for the discriminator. IC-GAN can generate realistic and diverse images for both labeled and unlabeled datasets, and can transfer to unseen datasets by changing the conditioning instances. IC-GAN can also control the semantics and style of the generated images by swapping the class labels or the instance features. However, using training data itself for retrieval potentially limits the generalization capacity.
KNN-Diffusion [149] uses large-scale retrieval methods to train a diffusion-based model without any text data. The model is conditioned on two inputs: a text or image embedding extracted by CLIP, and the k nearest neighbors of the embedding from a large image database. The kNN embeddings help to bridge the gap between the text and image distributions, and to generate images from different domains by simply swapping the database. The model also enables text-driven local semantic manipulations without masks, by fine-tuning the model to predict the original image from a manipulated version. RDM [150] combines a small diffusion or autoregressive model with a large external image database into a semi-parametric model. During training, the model retrieves a set of nearest neighbors from the database for each training image and conditions the generative model on their CLIP embeddings. This way, the model learns to compose new scenes based on the retrieved visual content. During inference, the model can generalize to novel domains, tasks, and conditions by changing the database or the retrieval strategy. Re-imagen [148] retrieved information from an external multi-modal knowledge base to produce realistic and faithful images, especially for rare or unseen entities. It is based on a cascaded diffusion model that conditions its generation on both the text prompt and the retrieved image-text pairs. It also proposed an interleaved classifier-free guidance schedule to balance the alignment between the text and the retrieval conditions. Re-imagen achieves state-of-the-art performance on COCO and WikiImages datasets, and significantly improves the fidelity of generated images on a new benchmark called EntityDrawBench. X&Fuse [233] is a general approach for conditioning on visual information when generating images from text. It works by concatenating the conditioned image and the noised image before each attention block in a U-Net architecture, and allowing interaction between them via self-attention. Retrieve&Fuse is a special case of X&Fuse, where the conditioned image is retrieved from a large bank of images using a text or image index. X&Fuse has several advantages over alternative methods, such as robustness to spatial differences, no information loss. RPG [77] retrieves representative images to construct informative in-context examples (i.e., image-region pairs), and utilizes multimodal chain-of-thought reasoning [234] to plan out complementary subregions for compositional text-to-image diffusion.
@@ IV-D2 Image Captioning
Retrieval-augmented Image Captioning is an innovative approach inspired by retrieval-augmented language generation. Unlike traditional image captioning methods that rely solely on the input image, this novel technique usually generates descriptive sentences by leveraging a set of captions retrieved from a datastore in addition to the image itself. Memory-augmented (MA) [162] exploits explicit knowledge from a memory bank to improve the caption generation. The memory bank is constructed by encoding the history context and the target word of each image-text pair in the training data, i.e., image features and past sentence (key) and target word (value). During inference, the model queries the memory bank with the current context vector and retrieves the most similar entries. Then, it computes a distribution over the vocabulary based on the retrieved entries and interpolates it with the original model’s prediction. The main highlights of MA are that it is non-parametric, compatible with any captioning model, and can adapt to larger datasets without additional training. RAMP [235] enhances image paragraph captioning by incorporating the best-retrieved candidate captions through adversarial training. It uses these captions as reference during the training of the discriminator, prompting the generator to include informative content from the retrieved captions into the generated caption. Additionally, RAMP employs a dynamic memory-augmented attention network to track coverage information and attention history, which helps in avoiding repetitive or incomplete descriptions. A copying mechanism is also integrated, allowing the selection of words from retrieved captions to improve the fluency and informativeness of the generated captions. RA-Transformer [44] adopts a kNN memory to retrieve relevant captions from an external corpus and augment the generation process. It consists of a Transformer-decoder architecture with a kNN-augmented attention layer that combines masked self-attention over the input tokens and cross-attention over the retrieved captions. A learned gate is used to balance the contribution of the local context and the external memory. For retrieval, it employs a CLIP-based visual encoder to perform approximate kNN searches in a visual-semantic space. RA-Transformer achieves competitive performance on the COCO dataset and demonstrates the effectiveness of retrieval abilities for image captioning. EXTRA [236] enhances image captioning by incorporating captions retrieved from similar images. It uses a pretrained Vision and Language (V&L) BERT encoder to jointly process the image and the retrieved captions, creating a rich cross-modal representation. The decoder then attends to both visual and linguistic representations, benefiting from the additional textual context. EXTRA shows that retrieving a sufficient number of captions (e.g., k=5) indeed improves captioning quality. REVEAL [237] is a retrieval-augmented visual language model that learns to encode and retrieve world knowledge from a large-scale memory, using a unified encoder and a diverse set of multimodal sources. It consists of four components: the memory, the encoder, the retriever, and the generator. For image captioning, REVEAL takes an image as input, retrieves relevant knowledge entries from the memory, and fuses them with the image features by introducing an attentive fusion layer that injects retrieval score into the attention calculation procedure. REVEAL is pre-trained on a massive image-text corpus with four knowledge sources: Wikipedia passages and images, web images with alt-text captions, knowledge graph triplets, and visual question answering pairs. SMALLCAP [45] generates a caption conditioned on an input image and related captions retrieved from a datastore of text via image-to-text retrieval. It uses a pre-trained CLIP vision encoder and GPT-2 language decoder, connected by trainable cross-attention layers. The retrieved captions are used as a prompt to the decoder, providing a task demonstration tailored to the input image. SMALLCAP differs from previous work in two main ways, i.e. SMALLCAP employs a straightforward prompt-based conditioning method, wherein retrieved captions are used as a prompt to a language model. Moreover, SMALLCAP is the first to leverage retrieval augmentation for training-free domain transfer and generalization in image captioning task. The Cross-Modal Retrieval and Semantic Refinement (CRSR) method [238] enhances remote sensing image captioning by introducing a novel approach that combines cross-modal retrieval with semantic refinement. It starts by employing a cross-modal retrieval model to fetch relevant sentences for each image, using these sentences as primary semantic information to support the captioning process. To refine this information, a semantic refinement module filters out misleading details and emphasizes visually salient semantic content. Additionally, a Transformer Mapper network is introduced to expand the representation of image features with learnable queries, capturing more intricate details within the images. The integration of refined semantic tokens and enriched visual features into a cross-modal decoder results in accurate and contextually rich captions.
@@ IV-E2 Video Generation
Text-to-video generation is to generate video given natural language descriptions. Animate-A-Story [187] develops a framework which can generate high-quality storytelling videos based on texts. It first separates the text into individual plots, and decorates the description using LLM. It then retrieves relevant videos for each plot through a dense retriever [109]. It generates videos through a latent diffusion model, consisting of two branches: a text encoder CLIP, and a structure encoder which takes the estimated depth of the retrieved videos as structure control.
@@ IV-F1 Text-to-3D
Retrieval can be applied to augment the generation of 3D contents. ReMoDiffuse [49] aims at generating motions using diffusion models. It first retrieves relevant motion entites through CLIP given text input, then leverages the information of the text and the retrieved entities through a semantic-modulated attention layer. AMD [156] designs two branches of motion diffusion for fidelity and diversity. The first branch inputs the original prompt text for diffusion; the second branch decomposes the input text into anatomical scripts and retrieve similar reference motions for diffusion. A transformer-based fusion module is further applied to adaptively balance the result from two branches. RetDream [48] targets general 3D generation, using retrieved 3D assets to augment the process of variational score distillation from 2D diffusion models. Given an input query, it retrieves relevant 3D assets through CLIP, then utilizes the retrieved assets to provide geometric prior and adapted 2D prior. Concretely, retrieved assets not only impose an additional velocity on particles for distribution initialization, but also help optimize the 2D diffusion model through low-rank adaptation.
@@ IV-G RAG for Knowledge
Structured knowledge, including knowledge base and knowledge graph, is widely used in language tasks. It usually serves as the retrieval source to augment generation.
@@ IV-G1 Knowledge Base Question Answering
Knowledge base question answering involves leveraging a knowledge base to identify the correct answer to a question. Many semantic parsing methods are proposed, which generate logical forms (e.g. SPARQL) given a question.

### v2

@@ IV-B4 Automatic Program Repair
Buggy code can take a lot of effort to fix. Automatic program repair leverages generative models to output the correct version. Query-based RAG techinque is widely used in automatic program repair [213, 131, 132, 187, 212]. Among them, RING [213] retrieves similar error messages based on both sparse and dense vectors, then builds prompt for the generator Codex [2]. CEDAR [131] applies for both assertion generation and program repairs tasks; it uses sparse and dense retrieval to search for similar codes, then forms prompt for Codex to generate results. InferFix [132] crafts a prompt carrying the bug type, location, relevant syntax hierarchies, and similar fixes through dense retrieval. Then it also uses Codex for generation. RAP-Gen [187] also retrieves similar buggy codes and corresponding fixes through hybrid retriever (including both sparse and dense retriever). It finetunes CodeT5 [106] with this RAG paradigm. SARGAM [212] searches similar buggy code using dense retrieval, generates patches using augmented input, then applies another model to modify the final result. These research works also involve error localization, which is not our focus. RTLFixer [285] leverages ReAct and RAG to implement an agent fixing errors in Verilog codes. It iteratively retrieves relevant errors and corresponding solutions, and combines reasoning and action planning to form prompts for LLMs.
@@ IV-B5 Text-to-SQL and Code-based Semantic Parsing
Semantic parsing is the task of translating natural language utterances to unambiguous structured meaning representations, where code language is often leveraged to augment this process. SQL is not only a programming language but can also be considered as a structured representation, so we place text-to-SQL (a special case of code generation) in this subsection. Related research works all apply query-based RAG. XRICL [194] focuses on the problem of translating non-English utterances into SQL queries. It searches and reranks English utterance using non-English ones by dense retrieval, then builds prompt for Codex to generate SQL queries. SYNCHROMESH [121] proposes constrained semantic decoding to enforce rich syntactic and semantic constraints during generation of SQL or other languages. It uses the similarity between abstract syntax trees (AST) to finetune the dense retriever S-Bert. During inference, similar NL and SQL are retrived to form the prompt of GPT-3. CodeICL [286] uses Python for semantic parsing, and augments prompts with a structured domain description for GPT generation. In few-shot learning setting, it leverages BM25 sparse retrieval to find similar training examples. RESDSQL [287] ranks schemas using cross-encoder, then includes ranked schemas into prompts to generate SQL skeleton and SQL query. ReFSQL [288] retrieves relevant questions and corresponding SQL to augment text-to-SQL generation. It involves structure-enhanced retriever with schema linking, and Mahalanobis contrastive learning to improve the retrieval performance. ODIS [289] retrieves in-domain and out-of-domain demonstrations using BM25, then includes them for in-context learning to generate SQL. Another work [290] retrieves both similar and diverse demonstrations, and then builds prompt for in-context learning to generate SQL. MURRE [291] conducts multi-hop retrieve-rewrite, where relevant tables are retrieved through dense retrieval and then re-writed to generate new tabularized question. A rank module at last integrate the retrieval results and select the top tables for the text-to-SQL module. CodeS [292] retrieves relevant information from table databases in a coarse-to-fine manner, then includes retrieved values to build prompts for both finetuning and inference.
@@ IV-B6 Others
There are several other code-related tasks that adopt RAG paradigm. In  [293] for numerical reasoning task, the Chain-of-Thought is replaced with the programs as the intermediate reasoning step, and dense retrieval-based similar examples are augmented in prompt for downstream LLMs. De-fine [294] tries to resolve intricate tasks using programs. It follows the paradigm in SKCODER, retrieves relevant pairs of query and code, then produces sketch template for real program generation. After generation, it combines the feedback to refine the answer with the same generator. The refined programs, regarded as optimal results, are added back to the codebase for subsequent serving. E&V [295] leverages an LLM-based agent for program static analysis. The agent uses source code retrieval, pseudo-code execution, execution specifications verification, and other tools to form intermediate results. The retrieval is implemented by AST pattern matching. Code4UIE [296] leverages code representation for information extraction tasks. It retrieves relevant examples through dense retrieval, and constructs in-context learning prompt using the retrieved queries and their corresponding codes. StackSpotAI [297] builds an AI coding assistant, which incorporates many code-related tasks. It implements an RAG component, identifying the most relevant pieces of information which serve as the context for GPT generation. InputBlaster [298] aims to generate unusual text input that could cause mobile app crash. It combines retrieved relevant buggy text with valid input to form the prompt for generation.
@@ IV-C1 Audio Generation
The goal of audio generation is to generate audio with natural language input.
Figure 11: Architecture of Re-AudioLDM [158] model.
Re-AudioLDM [158] adopts dense retriever CLAP [26] to retrieve similar caption-audio pairs given input prompt. As shown in Fig. 11, the generator, latent diffusion model and VAE-based decoder, take the representations of input text and retrieved pairs as input and generate output audio. Make-An-Audio [43] uses dense retriever CLAP [26] to augment data, retrieving related audio given natural language text. It then constructs pseudo prompts for diffusion-based text-to-audio model training.
@@ IV-D1 Image Generation
The retrieval process can not only help yield high-quality images even for rare or unseen subjects, but also reduces the parameter count and computational expense [44, 138, 150, 151, 151, 149, 303, 152, 153]. For GAN-based model, RetrieveGAN [44] employs a differentiable retrieval process to select compatible image patches for generation, with Gumbel-softmax trick and end-to-end training. IC-GAN [138] models data as a mix of conditional distributions around each training instance, conditioning both the generator and discriminator on these instances, and can control the semantics and style by swapping class labels or conditional instances. Recently, diffusion models beat GANs on image generation [304]. KNN-Diffusion [150] trains a text-to-image diffusion model without text data, by conditions the model on CLIP joint embedding of the instance and $k$-nearest neighbors from image database. These $k$-NN embeddings bridge the text-image distribution gap and allow image generation from different domains by swapping the database. Similarly, RDM [151] conditions diffusion or autoregressive models on CLIP embeddings of external image databases. It enables post-hoc conditioning on class labels, text prompts and zero-shot stylization [152]. Beyond retrieving only images, Re-imagen [149] conditions on both text prompts and retrieved image-text pairs for text-to-image generation. Interleaved classifier-free guidance is also proposed to balance the alignment between text prompts and retrieval conditions. To avoid information loss of CLIP embeddings and access to all visual condition, Retrieve&Fuse [303] concatenates the retrieved conditional image and the noised image before each attention block in U-Net, and allowing interaction via self-attention. RPG [78] retrieves representative images to construct informative in-context examples (i.e., image-region pairs), and utilizes multi-modal chain-of-thought reasoning [305] to plan out complementary subregions for compositional text-to-image diffusion.
@@ IV-D2 Image Captioning
Retrieval-augmented image captioning typically synthesises description with a collection of retrieved captions, instead depending only on the input image. MA [163] augments via a memory bank, built with historical context and target word of image-text training set, and queried during inference with the current context. Vocabulary distribution is computed based on retrieved entries, and interpolated with the original prediction. In adversarial training, RAMP [306] employs retrieved captions as reference for discriminator training, prompting the generator to make full use of retrieved captions. The memory-augmented attention and copying mechanism are also exploited to better use. The RA-Transformer [45] and EXTRA [307], both retrieval-augmented transformer-based captioning models, utilize cross-attention over encoded retrieved captions.
Figure 12: Architecture of EXTRA [307] model.
EXTRA, as depicted in Fig. 12, jointly process the image and retrieved captions with V&L encoder, such that the decoder attends to both visual and linguistic contexts. Beyond retrieved captions, REVEAL [308] uniformly encodes and retrieves multi-modal world knowledge, including image-text pairs, question answering pairs, and knowledge graph triplets, which is then integrated with image features by retrieval score-aware attention module. Straightforwardly, SMALLCAP [46] employs a CLIP vision encoder and a LLM decoder, linked by trainable cross-attention layers, where retrieved captions serve as input-specific in-context examples for prompt for enhancement. For remote sensing image, CRSR [309] enhances retrieved captions with semantic refinement, i.e. filters out misleading details and emphasizes visually salient content. Besides, the visual features is also enriched by transformer network with learnable queries, capturing more intricate details within the images
IV-D3 Others
There also exist many retrieval augmented works for other image-related tasks. For visual question answering (VQA), PICa [310] leverages GPT-3’s implicit knowledge retrieval and reasoning capabilities, which converts images into textual descriptions, and then prompts GPT-3 to predict answers based on these descriptions and the question. RA-VQA [311] identifies a limitation that the retrieval in previous work is trained separately from answer generation, and propose a joint training scheme that integrates differentiable retrieval with answer generation, enabling end-to-end training. For visually grounded dialogue, KNN-based Information Fetching (KIF) [312] enhances generative Transformer for dialog modeling. Each KIF module learns a read operation to access fixed external knowledge. Maria [313], a neural conversation agent, enhances dialog generation by leveraging visual experiences retrieved from a large-scale image index as extra context. For multi-modal machine translation, which aims to improve NMT with multi-modal information, [314] incorporates visual information at the phrase level to address the sparsity of paired sentence-image, employing a conditional VAE to filters out redundant visual information from sentence-image datasets.
@@ IV-E2 Video Generation
Text-to-video generation is to generate video given natural language descriptions.
Figure 13: Architecture of Animate-A-Story [210] model.
As shown in Fig. 13, Animate-A-Story [210] develops a framework which can generate high-quality storytelling videos based on texts. It first separates the text into individual plots, and decorates the description using LLM. It then retrieves relevant videos for each plot through a dense retriever [110]. It generates videos through a latent diffusion model, consisting of two branches: a text encoder CLIP, and a structure encoder which takes the estimated depth of the retrieved videos as structure control.
IV-E3 Others
There also exist many retrieval augmented works for other video-related tasks. For video question answering (VideoQA), MA-DRNN [315] leverages the differentiable neural computer (DNC) with an external memory, for storing and retrieving useful information in questions and videos, and modeling the long-term visual-textual dependence. Given the video input, R2A [316] retrieves semantically similar texts by multi-modal model, e.g. CLIP, and query LLM with both the question and the retrieved texts. For video-grounded dialogue, [317] proposes TVQA+ dataset which enables to retrieve relevant moments and visual concepts to answer questions about videos, and also proposes Spatio-Temporal Answerer with Grounded Evidence (STAGE) to exploit it. VGNMN [318] also extracts visual cues from videos, while the retrieval is carried out using neural module networks parameterized by entities and actions in previous dialogues. VidIL [319] exploits image-language models to translate video content into temporal-aware prompts with few-shot examples, for various video-language tasks including video captioning, video question answering, video caption retrieval, and video future event prediction. Notably, for trustworthy autonomous driving, RAG-Driver [320] grounds the MLLM in relevant expert demonstrations from a memory database, to produce driving action explanations, justifications, and control signal prediction.
@@ IV-F1 Text-to-3D
Retrieval can be applied to augment the generation of 3D contents. ReMoDiffuse [50] aims at generating motions using diffusion models. It first retrieves relevant motion entites through CLIP given text input, then leverages the information of the text and the retrieved entities through a semantic-modulated attention layer.
Figure 14: Architecture of AMD [157] model.
AMD [157] designs two branches of motion diffusion for fidelity and diversity. As shown in Fig. 14, the first branch inputs the original prompt text for diffusion; the second branch decomposes the input text into anatomical scripts and retrieve similar reference motions for diffusion. A transformer-based fusion module is further applied to adaptively balance the result from two branches. RetDream [49] targets general 3D generation, using retrieved 3D assets to augment the process of variational score distillation from 2D diffusion models. Given an input query, it retrieves relevant 3D assets through CLIP, then utilizes the retrieved assets to provide geometric prior and adapted 2D prior. Concretely, retrieved assets not only impose an additional velocity on particles for distribution initialization, but also help optimize the 2D diffusion model through low-rank adaptation.
@@ IV-G RAG for Knowledge
Structured knowledge, including knowledge graphs and tables, is widely used in language-related tasks. It usually serves as the retrieval source to augment generation. In addition to normal sparse retrieval and dense retrieval, NER (named-entity recognition) technique is applied to identify specific entites in queries, and graph-aware neighbor retrieval is leveraged to extract relevant entities and relations.
@@ IV-G1 Knowledge Base Question Answering
KBQA (knowledge base question answering) typically leverages a knowledge base to identify the correct answer to a question. Many semantic parsing methods are proposed, which generate logical forms (e.g. SPARQL) given a question.